What is the expression and answer to this problem?
Shakira has 3 2/5 pounds of jelly beans. She wanted to take half of the jelly beans and divide them amongst 6 of her friends. Write an expression that could be used to show How much each friend Should get.

Answer by stanbon(75887)

What is the expression and answer to this problem?
Shakira has 3 2/5 pounds of jelly beans. She wanted to take half of the jelly beans and divide them amongst 6 of her friends. Write an expression that could be used to show How much each friend Should get.
Ans: (1/6)[3 1/2]/2 = (1/6)(7/2)(1/2) = 7/24 lbs.
===============================================
Cheers,
Stan H.
---------------